OpenStudy (freemap):

Connor is using a computer program to graph the function J(n) = –13n – 421. He wants to see the intercepts. Which minimum values should he use to set up the viewing window? A. X Min: –40, Y Min: –500 B. X Min: –25, Y Min: –400 C. X Min: –400, Y Min: –25 D. X Min: –500, Y Min: –40

OpenStudy (freemap):

I think A?

OpenStudy (texaschic101):

J(n) = -13n - 421 y = -13n - 421 the y intercept is -421 the x intercept can be found by subbing in 0 for y 0 = -13n - 421 13n = -421 n = -421/13 n = -32 and 5/13 So I am getting x intercept as -32 5/13 and y int as -421....but I am not real sure of the answer...sorry

OpenStudy (texaschic101):

I am thinking B

OpenStudy (freemap):

Thank's for solving for me, appreciate that. I got 32.38 so I thinking A? If not that's not right its okay, thanks for helping

OpenStudy (freemap):

Just saw your post, So B? ok

OpenStudy (texaschic101):

because with negatives the smaller number is bigger

OpenStudy (freemap):

ok thanks, I have two more questions Do you mind helping? If you have something else to do I understand

OpenStudy (texaschic101):

I can try to help...can't make no promises that I will know the answer

OpenStudy (freemap):

Great, ok thanks